OpenStudy (anonymous):

MEDALS GIVEN What is the quotient of two over three divided by three over ten in simplest form? one over five five over thirteen two and two over nine three and one over three

OpenStudy (solomonzelman):

\[\frac{a}{b} \div \frac{x}{z}~~~~~~~~->~~~~~~~~~~~~\frac{a}{b} \times \frac{z}{x}\] (quotient is the result from division.)